OpenStudy (anonymous):

solve each equation or inequality. 4 log ( x+3)=9

OpenStudy (anonymous):

\[\log(x+3)=\frac{9}{4}\] use \[\log a=b \implies 10^b=a\]

OpenStudy (anonymous):

plug the numbers into the equation ?

OpenStudy (anonymous):

yes these are logarithmic laws

OpenStudy (anonymous):

\[10^{9/4}=x+3\]

OpenStudy (anonymous):

and that's my final answer?

OpenStudy (anonymous):

\[x=10^{9/4}-3\]

OpenStudy (anonymous):

that's my final step ?

OpenStudy (anonymous):

@Jonask

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you understand

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

yes , thanks .